# Event Camera Meets Resource-Aware Mobile Computing:

### v3

@@ 4.3. Event-based matching
Motivation. Matching involves identifying corresponding features between event streams captured at different times or viewpoints, forming the basis for tasks like visual odometry, video interpolation, and mobile sensing. Event cameras, unlike RGB cameras that operate at fixed intervals, trigger only on brightness changes, making them ideal for dynamic scenes. They capture edges and textures efficiently while avoiding redundant data. The precise temporal information in event data also enables accurate motion inference. By mitigating issues like motion blur and low frame rates common in traditional cameras, event cameras offer distinct advantages for matching tasks—even with single-modal data—demonstrating strong performance in high-frequency, fast-changing environments.
Challenge. Event-based matching faces major challenges due to the sparsity of events in pixel space and their uneven temporal distribution, which hinder robust feature extraction—especially in low-light or static scenes. This calls for specialized methods adapted to event camera properties. Moreover, the ultra-high temporal resolution produces massive data volumes, raising computational costs. Achieving real-time performance in resource-limited settings thus requires algorithms that balance efficiency and accuracy, minimizing latency while ensuring reliability in dynamic env..
@@
Local feature matching [131, 132, 133, 134, 135, 136, 137, 138, 139, 140] Optical Flow Feature descriptor: SIFT/SURF/ORB Batch Low latency, noise resistance Low computational resource Difficult to handle global motion Difficult to handle repeat texture Need further extraction method
Optimization-based matching [141, 142, 143] Contrast maximization Levenberg-Marquardt Graph optimization Batch/Asynchronous Global consistency Higher accuracy after iterations Detailed design of optimizer Higher computational complexity Higher time latency Hard to converge Sensitive to initial values
Neural network-based matching [144, 145, 146, 147, 148] CNN Batch/Asynchronous Adaptive learning Less manual intervention Efficient inference Large amounts of training data Need parameter tuning Poor generalization ability
@@
Local feature matching methods. These methods focus on detecting and associating local features within visual data [149]. Techniques such as optical flow [131, 132] and descriptors like SIFT [135], SURF [136], and ORB [137] exemplify this approach. They offer low latency, robustness to noise, and low computational demands, making them well-suited for mobile, autonomous, and robotic applications [139, 140]. However, local feature matching faces challenges with event streams, especially under large global motion, repetitive textures, or heavy reliance on specific feature extractors [150]. Its performance also declines in scenes with sparse features or dramatic changes, underscoring the need for more advanced and robust techniques.
Optimization-based matching methods. These methods aim to ensure global consistency by solving optimization problems to align event data or trajectories. Techniques such as contrast maximization [141, 142, 143], the Levenberg-Marquardt algorithm, and graph-based optimization are widely used. They offer high flexibility and accuracy, supporting both batch and asynchronous processing, which makes them valuable in event camera applications. However, these approaches are computationally demanding and sensitive to initial conditions—poor initialization can lead to local optima or non-convergence. Their high memory and processing requirements also limit real-time usability, especially for large-scale data. Achieving a balance between efficiency and accuracy remains a core challenge.
Deep learning-based matching methods. These methods mark a transformative shift by employing CNNs [146], SNNs [144, 145], and GNNs [148] to automatically learn representations and matching strategies, surpassing handcrafted descriptors in accuracy. GNNs are well-suited to the asynchronous nature of event streams and excel in high-speed, dynamic environments. However, these models are computationally demanding, rely on large annotated datasets, and struggle to generalize, while training remains time-consuming and requires careful hyperparameter tuning.
@@ 4.4. Event-based mapping
Motivation Building on event matching, asynchronous event streams from different viewpoints can be fused across poses to incrementally reconstruct a dense 3D map. Event-based sensing leverages spatial sparsity and high temporal resolution, reducing computation by focusing on regions of change and minimizing motion blur to preserve edges during fast motion. These properties enable real-time, efficient, and robust mapping in dynamic environments.
@@
Types of Mapping Methods Advantages Disadvantages or Challenges Frame-based mapping[151, 152, 153, 154, 155, 156] ✓ Improves the accuracy of depth estimation ✓ Enhances the quality of map construction ✓ Lower latency and computation cost ✗ Requires measuring and updating depth information between current and initial detection frames, potentially increasing computational load Filter-based mapping [157, 158, 159, 160] ✓ High robustness, suitable for rapidly changing environments ✓ Dynamically adjusts the system’s map representation ✗ Requires a large number of parameters to represent camera poses, potentially increasing computational complexity    Continuous-time mapping [161, 162, 163] ✓ Reduces parameter complexity ✓ Improves mapping accuracy and efficiency ✓ Simultaneously updates camera poses and 3D landmarks ✗ Requires handling continuous curve interpolation and optimization problems ✗ Increasing computational difficulty ✗ High latency due to frequent states update Spatio-temporal Consistency[164] ✓ Improves the accuracy of map construction ✓ Optimizes motion parameters ✗ Requires iteratively searching for the closest points and applying a pruned ICP algorithm ✗ High computational cost and latency
@@
Frame-based mapping methods. Frame-based mapping methods often use event-derived 2D representations with depth filters to iteratively refine scene depth via feature triangulation. For example, [153] estimates poses relative to planar structures by minimizing reprojection errors, while [155] adopts the SVO algorithm [156] for pose estimation from event feature correspondences. These methods typically model depth with Gaussian–uniform filters, updating estimates from feature comparisons guided by poses. Additionally, [152] improves mapping by solving pose estimation as a least-squares problem on 2D–3D line constraints.
Filter-based mapping methods. Filter-based mapping techniques address the asynchronous nature of event data by continuously updating maps during camera tracking. Line-based vSLAM refines maps by measuring distances between incoming events and reprojected 3D lines, generating point cloud reconstructions. These methods often use the Hough transform to extract 3D line features, leveraging spatial correlations for robustness in dynamic scenes. For example, [157] estimates poses via distances between back-projected event rays and planar points, while [160] improves precision with a probabilistic measurement function on planar surfaces. More recent strategies, such as [158], update filter states by evaluating event-to-line distances, often combining EMVS [159] with Hough-based line extraction to strengthen event-line associations and enhance spatial accuracy.
Continuous-Time mapping methods. To reduce the high parameter count in filter-based discrete pose representations, continuous-time mapping replaces discrete poses with smooth trajectory models such as B-splines or Gaussian processes, enabling interpolation from local control states. This lowers complexity and supports joint optimization of poses and landmarks, improving accuracy and efficiency. For example, [161] employs B-splines, while [162] use Gaussian process motion models to interpolate poses at arbitrary timestamps. [162] further integrates incremental SfM for consistent refinement of control states and landmarks, and [164] introduces a spatio-temporal constraint based on equal-time event pairs to enhance rotational accuracy. In practice, these methods iteratively establish correspondences and apply pruned ICP for spatial consistency, yielding more precise event-based maps.
@@ 5. Acceleration
Event-based sensing generates data only when pixel intensity changes, drastically reducing redundancy and offering efficiency, low power, and minimal latency, which is well-suited for high-speed, energy-constrained applications. With growing demand on mobile and edge platforms such as drones, vehicles, and wearables, real-time and efficient processing is critical but remains challenging under limited resources. Balancing accuracy and efficiency thus becomes central for deployment. To address this, optimized hardware accelerators and specialized software must exploit event sparsity to cut computation while preserving accuracy (Fig.7, Fig.8), enabling real-time event-based vision on mobile devices for advanced autonomous applications.
@@ 5.1. Hardware acceleration
Motivation. Event-based vision transforms visual processing by generating data only from scene changes, achieving sparsity, high temporal resolution, and low latency—ideal for real-time, power-limited applications. Yet, conventional processors built for dense synchronous data struggle with sparse, asynchronous streams, making specialized hardware essential. Hardware accelerators bridge raw events to high-level tasks like detection, tracking, and reconstruction by exploiting sparsity and parallelism to cut redundancy, streamline data flow, and support massive parallel processing. This enables precise, low-latency, and energy-efficient event-based systems ready for practical deployment.
@@
Literature review. Recent research on hardware acceleration for event-based vision can be broadly categorized into three main approaches: neuromorphic computing, event-driven deep neural network (DNN) acceleration, and hardware optimization techniques aimed at enhancing efficiency and reducing power consumption.
Neuromorphic computing for event-based vision. Neuromorphic computing, inspired by the brain’s dynamic processing, mimics biological neurons and synapses to enable efficient event-driven computation. SNNs typically run on custom neuromorphic chips for low-power, high-efficiency processing. Notable devices include BrainScales [165], Spinnaker [166], Neurogrid [167], TrueNorth [168], Darwin [169], and more recently, Loihi [170], Tianjic [171], and Speck [172]. A key challenge is improving energy efficiency via high-level brain-inspired mechanisms. Among these, the asynchronous chip Speck [172] stands out as a sensing-computing SoC that fully leverages sparse, event-driven processing. Operating at ultra-low power (0.70 mW in real-time), Speck demonstrates neuromorphic computing’s promise for power-constrained mobile and edge systems.
Event-driven DNN acceleration. Neuromorphic computing provides a bio-inspired paradigm for event-driven processing, while another crucial research direction focuses on hardware acceleration tailored to event-based deep learning models. Unlike conventional GPUs, which fail to fully exploit the sparsity of event data, specialized architectures have been developed to achieve higher efficiency. $(i)$ Sparse dataflow architectures: The combinable dynamic sparse dataflow architecture (ESDA) [173] realizes a configurable sparse dataflow model on FPGAs. By employing a unified sparse token–feature interface to interconnect parameterizable network modules, ESDA reduces both latency and power consumption during event-based DNN inference, making it well-suited for edge deployment. $(ii)$ Optimized fusion for event-based vision: EventBoost [174] accelerates event–image fusion through a dedicated hardware accelerator on the Zynq SoC platform. By optimizing fusion workloads in real time, EventBoost mitigates inefficiencies typical of CPU- and GPU-based processing, significantly boosting performance for event-driven visual tasks.
@@
Hardware optimization for efficient event processing. Beyond neuromorphic computing and event-driven DNNs, hardware acceleration is essential for unlocking the full potential of event-based vision, particularly in power-constrained and efficiency-critical scenarios. One optimization strategy aims to reduce power consumption directly at the sensor interface. For example, [182] proposes a dedicated on-chip DVS interface that aggregates asynchronous event streams into ternary event frames, substantially lowering the energy demands of subsequent processing stages.
@@
In addition to such front-end innovations, a variety of FPGA-based accelerators have been explored to enhance efficiency across the entire event-vision pipeline by exploiting the inherent parallelism of event data. For instance, [176] presents a reconfigurable processing element architecture that integrates both a median filter core and an AI accelerator core for CNN inference within a system-on-chip (SoC). Leveraging Reconfigurable Multiple Constant Multiplication (RMCM) for efficient resource sharing, this design achieves an energy consumption of only 593.4 nJ per inference under 65 nm technology, significantly reducing computational cost. Similarly, [183] develops a complete event-driven optical flow camera system with FPGA-based acceleration for key modules such as event-driven corner detection and adaptive block matching optical flow. Expanding toward heterogeneous architectures, [184] demonstrates an FPGA/ARM platform for Event-based Monocular Multi-View Stereo (EMVS), where algorithm restructuring and mixed-precision quantization boost throughput while minimizing memory footprint. Finally, [174] proposes a Zynq SoC–based event–vision fusion accelerator that employs hardware–software co-design to distribute tasks between FPGA and CPU, maximizing parallelism and efficiency in computationally demanding fusion algorithms.
@@
Application-specific hardware acceleration design. The development of specialized hardware architectures tailored to accelerate sThe development of specialized hardware architectures tailored to specific tasks has become a central focus in event-based vision research. By optimizing hardware resources for domain applications, these solutions achieve substantial performance gains. For instance, [176] introduces an energy-efficient processor-on-chip for hand gesture recognition. A key driver is aerial robotics, where stringent constraints on size, weight, and power demand highly optimized systems. [39] presents FPGA-based hardware–software co-designs that exploit FPGA reconfigurability to implement efficient pipelines for drone tasks such as navigation and obstacle avoidance. Complementarily, [185] demonstrates a neuromorphic approach using Intel Loihi, where event data feed directly into SNNs for end-to-end drone control, achieving high responsiveness with low power. These directions highlight the synergy between specialized hardware and event-based vision. As edge applications—such as autonomous driving, robotics, and AR—demand low-power, real-time performance, domain-optimized accelerators will be pivotal for practical deployment.

### v4

@@ 4.3. Event-based matching
Motivation. Matching involves identifying corresponding features between event streams captured at different times or viewpoints, forming the basis for tasks like visual odometry, video interpolation, and mobile perception. Event cameras, unlike RGB cameras that operate at fixed intervals, trigger only on brightness changes, making them ideal for dynamic scenes. They capture edges and textures efficiently while avoiding redundant data. The precise temporal information in event data also enables accurate motion inference. By mitigating issues like motion blur and low frame rates common in traditional cameras, event cameras offer distinct advantages for matching tasks, even with single-modal data, demonstrating strong performance in high-frequency, fast-changing environments.
Challenge. Event-based matching faces major challenges due to the sparsity of events in pixel space and their uneven temporal distribution, which hinder robust feature extraction, especially in low-light or static scenes. This calls for specialized methods adapted to event camera properties. Moreover, the ultra-high temporal resolution produces massive data volumes, raising computational costs. Achieving real-time performance in resource-limited settings thus requires algorithms that balance efficiency and accuracy, minimizing latency while ensuring reliability in dynamic env..
@@
Local feature matching [Li_2024_CVPR, ren2024simpleeffectivepointbasednetwork, 10342794, hou2022fefusionvprattentionbasedmultiscalenetwork, calonder2008keypoint, BAY2008346, 6126544, calonder2010brief, 10629077, wang2024fedetrkeypointdetectiontracking] Optical Flow Feature descriptor: SIFT/SURF/ORB Batch Low latency, noise resistance Low computational resource Difficult to handle global motion Difficult to handle repeat texture Need further extraction method
Optimization-based matching [Gallego_2018, 9156854, Stoffregen2019EventCC] Contrast maximization Levenberg-Marquardt Graph optimization Batch/Asynchronous Global consistency Higher accuracy after iterations Detailed design of optimizer Higher computational complexity Higher time latency Hard to converge Sensitive to initial values
Neural network-based matching [wu2024lightweighteventbasedopticalflow, chen2024eventbasedmotionmagnification, deng2022voxel, rebecq2019events, schaefer2022aegnnasynchronouseventbasedgraph] CNN Batch/Asynchronous Adaptive learning Less manual intervention Efficient inference Large amounts of training data Need parameter tuning Poor generalization ability
@@
Local feature matching methods. These methods focus on detecting and associating local features within visual data [10.1177/0278364914554813]. Techniques such as optical flow [Li_2024_CVPR, ren2024simpleeffectivepointbasednetwork] and descriptors like SIFT [calonder2008keypoint], SURF [BAY2008346], and ORB [6126544] exemplify this approach. They offer low latency, robustness to noise, and low computational demands, making them well-suited for mobile, autonomous, and robotic applications [10629077, wang2024fedetrkeypointdetectiontracking]. However, local feature matching faces challenges with event streams, especially under large global motion, repetitive textures, or heavy reliance on specific feature extractors [Whelan2015ElasticFusionDS]. Its performance also declines in scenes with sparse features or dramatic changes, underscoring the need for more advanced and robust techniques.
Optimization-based matching methods. These methods aim to ensure global consistency by solving optimization problems to align event data or trajectories. Techniques such as contrast maximization [Gallego_2018, 9156854, Stoffregen2019EventCC], the Levenberg-Marquardt algorithm, and graph-based optimization are widely used. They offer high flexibility and accuracy, supporting both batch and asynchronous processing, which makes them valuable in event camera applications. However, these approaches are computationally demanding and sensitive to initial conditions—poor initialization can lead to local optima or non-convergence. Their high memory and processing requirements also limit real-time usability, especially for large-scale data. Achieving a balance between efficiency and accuracy remains a core challenge.
Deep learning-based matching methods. These methods mark a transformative shift by employing CNNs [deng2022voxel], SNNs [wu2024lightweighteventbasedopticalflow, chen2024eventbasedmotionmagnification], and GNNs [schaefer2022aegnnasynchronouseventbasedgraph] to automatically learn representations and matching strategies, surpassing handcrafted descriptors in accuracy. GNNs are well-suited to the asynchronous nature of event streams and excel in high-speed, dynamic environments. However, these models are computationally demanding, rely on large annotated datasets, and struggle to generalize, while training remains time-consuming and requires careful hyperparameter tuning.
@@ 4.4. Event-based mapping
Motivation Building on event matching, asynchronous event streams from different viewpoints can be fused across poses to incrementally reconstruct a dense 3D map. Event-based perception leverages spatial sparsity and high temporal resolution, reducing computation by focusing on regions of change and minimizing motion blur to preserve edges during fast motion. These properties enable real-time, efficient, and robust mapping in dynamic environments.
@@
Types of Mapping Methods Advantages Disadvantages or Challenges Frame-based mapping[Mueggler2014Eventbased6P, Yuan2016FastLA, Bertrand2020EmbeddedEV, Kueng2016LowlatencyVO, Zhu2019NeuromorphicVO, Forster2014SVOFS] ✓ Improves the accuracy of depth estimation ✓ Enhances the quality of map construction ✓ Lower latency and computation cost ✗ Requires measuring and updating depth information between current and initial detection frames, potentially increasing computational load Filter-based mapping [Weikersdorfer2012EventbasedPF, Chamorro2022EventBasedLS, Rebecq2018EMVSEM, Weikersdorfer2013SimultaneousLA] ✓ High robustness, suitable for rapidly changing environments ✓ Dynamically adjusts the system’s map representation ✗ Requires a large number of parameters to represent camera poses, potentially increasing computational complexity Continuous-time mapping [Mueggler2015ContinuousTimeTE, Liu2022AsynchronousOF, Gentil2020IDOLAF] ✓ Reduces parameter complexity ✓ Improves mapping accuracy and efficiency ✓ Simultaneously updates camera poses and 3D landmarks ✗ Requires handling continuous curve interpolation and optimization problems ✗ Increasing computational difficulty ✗ High latency due to frequent states update Spatio-temporal Consistency[Liu2021SpatiotemporalRF] ✓ Improves the accuracy of map construction ✓ Optimizes motion parameters ✗ Requires iteratively searching for the closest points and applying a pruned ICP algorithm ✗ High computational cost and latency
@@
Frame-based mapping methods. Frame-based mapping methods often use event-derived 2D representations with depth filters to iteratively refine scene depth via feature triangulation. For example, [Bertrand2020EmbeddedEV] estimates poses relative to planar structures by minimizing reprojection errors, while [Zhu2019NeuromorphicVO] adopts the SVO algorithm [Forster2014SVOFS] for pose estimation from event feature correspondences. These methods typically model depth with Gaussian–uniform filters, updating estimates from feature comparisons guided by poses. Additionally, [Yuan2016FastLA] improves mapping by solving pose estimation as a least-squares problem on 2D–3D line constraints.
Filter-based mapping methods. Filter-based mapping techniques address the asynchronous nature of event data by continuously updating maps during camera tracking. Line-based vSLAM refines maps by measuring distances between incoming events and reprojected 3D lines, generating point cloud reconstructions. These methods often use the Hough transform to extract 3D line features, leveraging spatial correlations for robustness in dynamic scenes. For example, [Weikersdorfer2012EventbasedPF] estimates poses via distances between back-projected event rays and planar points, while [Weikersdorfer2013SimultaneousLA] improves precision with a probabilistic measurement function on planar surfaces. More recent strategies, such as [Chamorro2022EventBasedLS], update filter states by evaluating event-to-line distances, often combining EMVS [Rebecq2018EMVSEM] with Hough-based line extraction to strengthen event-line associations and enhance spatial accuracy.
Continuous-Time mapping methods. To reduce the high parameter count in filter-based discrete pose representations, continuous-time mapping replaces discrete poses with smooth trajectory models such as B-splines or Gaussian processes, enabling interpolation from local control states. This lowers complexity and supports joint optimization of poses and landmarks, improving accuracy and efficiency. For example, [Mueggler2015ContinuousTimeTE] employs B-splines, while [Liu2022AsynchronousOF] use Gaussian process motion models to interpolate poses at arbitrary timestamps. [Liu2022AsynchronousOF] further integrates incremental SfM for consistent refinement of control states and landmarks, and [Liu2021SpatiotemporalRF] introduces a spatio-temporal constraint based on equal-time event pairs to enhance rotational accuracy. In practice, these methods iteratively establish correspondences and apply pruned ICP for spatial consistency, yielding more precise event-based maps.
@@ 5. Acceleration
Event-based perception generates data only when pixel intensity changes, drastically reducing redundancy and offering efficiency, low power, and minimal latency, which is well-suited for high-speed, energy-constrained applications. With growing demand on mobile agents such as drones, vehicles, and wearables, real-time and efficient processing is critical but remains challenging under limited resources. Balancing accuracy and efficiency thus becomes central for deployment. To address this, optimized hardware accelerators and specialized software must exploit event sparsity to cut computation while preserving accuracy (Fig.7, Fig.8), enabling real-time event-based vision on mobile agents for advanced autonomous applications.
@@ 5.1. Hardware acceleration
Motivation. Event-based vision transforms visual processing by generating data only from scene changes, achieving sparsity, high temporal resolution, and low latency, which is ideal for real-time, power-limited applications. Yet, conventional processors built for dense synchronous data struggle with sparse, asynchronous streams, making specialized hardware essential. Hardware accelerators bridge raw events to high-level tasks like detection, tracking, and reconstruction by exploiting sparsity and parallelism to cut redundancy, streamline data flow, and support massive parallel processing. This enables precise, low-latency, and energy-efficient event-based systems ready for practical deployment.
@@
Literature review. Recent research on hardware acceleration for event-based vision can be broadly categorized into three main approaches: neuromorphic computing, event-driven deep neural network (DNN) acceleration, and hardware optimization techniques aimed at enhancing efficiency and reducing power consumption (Fig.7, Fig.8).
Neuromorphic computing for event-based vision. Neuromorphic computing, inspired by the brain’s dynamic processing, mimics biological neurons and synapses to enable efficient event-driven computation. SNNs typically run on custom neuromorphic chips for low-power, high-efficiency processing. Notable devices include BrainScales [5536970], Spinnaker [6515159], Neurogrid [6805187], TrueNorth [doi:10.1126/science.1254642], Darwin [MA201743], and more recently, Loihi [8259423], Tianjic [Pei2019Towards], and Speck [Yao2024Spike-based]. A key challenge is improving energy efficiency via high-level brain-inspired mechanisms. Among these, the asynchronous chip Speck [Yao2024Spike-based] stands out as a sensing-computing SoC that fully leverages sparse, event-driven processing. Operating at ultra-low power (0.70 mW in real-time), Speck demonstrates neuromorphic computing’s promise for power-constrained mobile and edge systems.
Event-driven DNN acceleration. Neuromorphic computing provides a bio-inspired paradigm for event-driven processing, while another crucial research direction focuses on hardware acceleration tailored to event-based deep learning models. Unlike conventional GPUs, which fail to fully exploit the sparsity of event data, specialized architectures have been developed to achieve higher efficiency. $(i)$ Sparse dataflow architectures: The combinable dynamic sparse dataflow architecture (ESDA) [Gao2024A] realizes a configurable sparse dataflow model on FPGAs. By employing a unified sparse token–feature interface to interconnect parameterizable network modules, ESDA reduces both latency and power consumption during event-based DNN inference, making it well-suited for edge deployment. $(ii)$ Optimized fusion for event-based vision: EventBoost [Cao2024EventBoost] accelerates event–image fusion through a dedicated hardware accelerator on the Zynq SoC platform. By optimizing fusion workloads in real time, EventBoost mitigates inefficiencies typical of CPU- and GPU-based processing, significantly boosting performance for event-driven visual tasks.
@@
Hardware optimization for efficient event processing. Beyond neuromorphic computing and event-driven DNNs, hardware acceleration is essential for unlocking the full potential of event-based vision, particularly in power-constrained and efficiency-critical scenarios. One optimization strategy aims to reduce power consumption directly at the sensor interface. For example, [RUTISHAUSER2023717] proposes a dedicated on-chip DVS interface that aggregates asynchronous event streams into ternary event frames, substantially lowering the energy demands of subsequent processing stages.
@@
In addition to such front-end innovations, a variety of FPGA-based accelerators have been explored to enhance efficiency across the entire event-vision pipeline by exploiting the inherent parallelism of event data. For instance, [Fu2024A] presents a reconfigurable processing element architecture that integrates both a median filter core and an AI accelerator core for CNN inference within a system-on-chip (SoC). Leveraging Reconfigurable Multiple Constant Multiplication (RMCM) for efficient resource sharing, this design achieves an energy consumption of only 593.4 nJ per inference under 65 nm technology, significantly reducing computational cost. Similarly, [9727106] develops a complete event-driven optical flow camera system with FPGA-based acceleration for key modules such as event-driven corner detection and adaptive block matching optical flow. Expanding toward heterogeneous architectures, [li2022eventorefficienteventbasedmonocular] demonstrates an FPGA/ARM platform for Event-based Monocular Multi-View Stereo (EMVS), where algorithm restructuring and mixed-precision quantization boost throughput while minimizing memory footprint. Finally, [Cao2024EventBoost] proposes a Zynq SoC–based event–vision fusion accelerator that employs hardware–software co-design to distribute tasks between FPGA and CPU, maximizing parallelism and efficiency in computationally demanding fusion algorithms.
@@
Application-specific hardware acceleration design. The development of specialized hardware architectures tailored to accelerate specific tasks has become a central focus in event-based vision research. By optimizing hardware resources for domain applications, these solutions achieve substantial performance gains. For instance, [Fu2024A] introduces an energy-efficient processor-on-chip for hand gesture recognition. A key driver is aerial robotics, where stringent constraints on size, weight, and power demand highly optimized systems. [xu2023taming] presents FPGA-based hardware–software co-designs that exploit FPGA reconfigurability to implement efficient pipelines for drone tasks such as navigation and obstacle avoidance. Complementarily, [9560881] demonstrates a neuromorphic approach using Intel Loihi, where event data feed directly into SNNs for end-to-end drone control, achieving high responsiveness with low power. These directions highlight the synergy between specialized hardware and event-based vision. As edge applications—such as autonomous driving, robotics, and AR—demand low-power, real-time performance, domain-optimized accelerators will be pivotal for practical deployment.